OpenStudy (anonymous):

Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations.

OpenStudy (anonymous):

Equation # y - 4 = 1(x - 0)

OpenStudy (anonymous):

Equation #2 y - 3 = 1(x - 1)

jimthompson5910 (jim_thompson5910):

Using the point slope formula y - y1 = m(x-x1) only equation 1 is correct since m = 1, x1 = 0 and y1 = 4

OpenStudy (saifoo.khan):

Eq1 is correct.

OpenStudy (saifoo.khan):

cz in second one, the x cordinate is incorrect.